OpenStudy (anonymous):

I'm about to cry, could someone please help me? The graph of f(x) = 2x + 4 shifts five units to the right when it is replaced with the graph of f(x) = 2x - k. What is the value of k? A.) -5 B.) -1 C.) 1 D.) 5

OpenStudy (anonymous):

@PixieDust1 could you please help me?

OpenStudy (pixiedust1):

I'd love to help you, friend, but I don't think I can on this one. I'm not certain on the answer, and like to be certain before I help out. Sorry friend. :(

OpenStudy (anonymous):

ok

OpenStudy (pixiedust1):

@RadEn @ranga @dumbcow @whpalmer4 @wolfe8

OpenStudy (pixiedust1):

Here I got someone who could help you @Loser66

OpenStudy (anonymous):

thank you

OpenStudy (phi):

Are you sure about the question? If you move the line 5 steps to the right (see file) the equation is y= 2x - 6 and 6 is not one of your choices.

OpenStudy (anonymous):

I'm absolutely positive about the question. I don't understand it either. :/

OpenStudy (phi):

moving the line to the right 5 steps makes sense, right? if you have a point on the (for example (0,4) is on the line), that point will be 5 steps to the right on the new line... at (5,4) ... you add 5 to the x value) the new line has the same slope, so you can think of it as point-slope problem slope is 2, point is (5,4) y-4 = 2(x - 5) in slope-intercept form you get y - 4 = 2x - 10 y = 2x - 6

OpenStudy (anonymous):

So, after looking at this problem, I can safely say that the answer is D., which is 5.

OpenStudy (phi):

the answer is 6 I would talk to your teacher about this one.

OpenStudy (anonymous):

Do you have time to help me w/ another algebra question?

OpenStudy (anonymous):

6?

OpenStudy (loser66):

Question: let say f(x1) = 2x+4 f(x2) = 2x-k which one is replaced one? which one is original graph?

OpenStudy (pixiedust1):

This happened to me once on a quiz. I gave my entire explanation of why none of the options work and what the answer should of been, and guess what? It was a trick question! I got bonus points :D

OpenStudy (anonymous):

The original one is f(x1) = 2x + 4. Probably wrong.

OpenStudy (anonymous):

I just need a walkthrough of the question. :-) I guess the answer, if it's wrong, tell me, if it's right just put a smiley face.

OpenStudy (anonymous):

Probably wrong though. :(

OpenStudy (loser66):

yea, me too. think of this, f(x) = 2x -1 then move it to the right 5 units, we get f(x) =2x+4

OpenStudy (loser66):

cross the fingers, hehehe

OpenStudy (loser66):

I think, it's b

OpenStudy (loser66):

oh, sorry, I think it's C , k = 1 because if k =1, f(x) = 2x -1 move to the right 5 units, f(x) = 2x -1 +5 = 2x +4

OpenStudy (anonymous):

So, I can safely say, the answer is C., which is 1.